OpenStudy (kendricklamar2014):

What is the slope of the line with equation 4x + 2y = 8?

OpenStudy (kendricklamar2014):

A. -2 B. -0.5 C. 0.5 D. 2

OpenStudy (solomonzelman):

subtract 4x from both sides, divide each term in the equation by 2. what do you get after this?

OpenStudy (kendricklamar2014):

So it would be like this: 4x - 4x + 2y = 8 -4x

OpenStudy (alexandervonhumboldt2):

no need in thats ax+by=c -a/b is the slope

OpenStudy (kendricklamar2014):

8 - 4x = ? @SolomonZelman

OpenStudy (solomonzelman):

yes, you would be getting, 2y=-4x+8. then divide each term by 2.

OpenStudy (solomonzelman):

AlexandervonHumboldt2 , it is easier visually to do it in a slope intercept form.

OpenStudy (kendricklamar2014):

2y / 2 = -4x + 8/2

OpenStudy (alexandervonhumboldt2):

i know that

OpenStudy (kendricklamar2014):

= y = -4x + 4

OpenStudy (solomonzelman):

2y/2 = -4x/2 + 8/2 like this

OpenStudy (solomonzelman):

you divide every term by 2.

OpenStudy (solomonzelman):

so what does THIS simplify to ?

OpenStudy (kendricklamar2014):

Im not sure??

OpenStudy (solomonzelman):

take a try

OpenStudy (kendricklamar2014):

y = 0.5??

OpenStudy (solomonzelman):

2y divided by 2 is ?

OpenStudy (solomonzelman):

y.

OpenStudy (solomonzelman):

-4x divided by 2 is? 8 divided by 2 is ?

OpenStudy (kendricklamar2014):

-4x / 2 = -2x

OpenStudy (kendricklamar2014):

8 / 2 = 4

OpenStudy (solomonzelman):

yes

OpenStudy (kendricklamar2014):

So now divide both sides by -2 right

OpenStudy (solomonzelman):

you already have: 2y+4x=8 2y+4x-4x=8-4x 2y=-4x+8 dividing each term by 2: y=-2x+4 (I write it altogether, after we have divided each term separately)

OpenStudy (solomonzelman):

in a slope intercept form, y=mx+b , the slope is m. You have: y=-2x+8. What is the slope?

OpenStudy (kendricklamar2014):

-2

OpenStudy (solomonzelman):

yes

OpenStudy (kendricklamar2014):

Thanks